Can you please solve this problem I have hard time to form the equation.
A ladder is resting against a wall. The top of the ladder touches the wall at a height of 6 feet. Find the length of the ladder if the length is 2 feet more than its distance from the wall.
Thanks

Answer by wuwei96815(245)

The hardest part of word problems is translating them into math problems. This problem is a right triangle problem. The formula for a right triangle is
a2 + b2 = c2. Let's assign values to these letters and see if we can't solve your problem:

Let a=6'
Let b=x because it is unknown
Let c=x+2 because it is 2' longer than b

a2+b2=c2 becomes (6)2 + (x)2 = (x+2)2

Let's solve for x:
36 + x2 = x2 +4x +4
36 - 4 = -x2 + x2 + 4x +4
32 = 4x
8 = x

If x=8 then x+2=10

If you are so inclined go back to the equation a2+b2=c2 and see if our answers are correct.